The senior classes at High School A and High School B planned separate trips to New York City. The senior class at High School A rented and filled 1 van and 6 buses with 372 students. High School B rented and filled 4 vans and 12 buses with 780 students. Each can and each bus carried the same number of students. How many students can a van carry? How many students can a bus carry?
Answer by ikleyn(52781)

.
The senior classes at High School A and High School B planned separate trips to New York City. The senior class at High School A rented and filled 1 van and 6 buses with 372 students. High School B rented and filled 4 vans and 12 buses with 780 students. Each can and each bus carried the same number of students. How many students can a van carry? How many students can a bus carry?
-------------------------------------------------

Your equations are

1*v +  6*b = 372,    (1)
4*v + 12*b = 780,    (2)

where v is a van capacity and b is a bus capacity.

Now express v = 372 - 6b from (1) and substitute it into (2). In this way you exclude v and obtain a single equation for b:

4*(372-6b) + 12b = 780.

Can you complete the solution on your own?